OpenStudy (iwanttogotostanford):

NEED HELP ASAP PLEASE: Solve x^2 + 8x − 3 = 0 using the completing-the-square method. x = four plus or minus the square root of three x = negative four plus or minus the square root of three x = four plus or minus the square root of nineteen x = negative four plus or minus the square root of nineteen

OpenStudy (iwanttogotostanford):

@pooja195 @Nnesha

OpenStudy (jack1):

complete the square is: turn this \(\large ax^2 + bx + c = 0\) into this \(\large a(x+d)2 + e = 0\) remember: \(\Large d = \frac b{2a} \) \(\Large e = c - \frac{b^2}{4a}\)

OpenStudy (iwanttogotostanford):

so it would be C right???

OpenStudy (jack1):

your equation is: \(\large x^2 + 8x + (-3) = 0\) so turn it into this: \(\large a(x+d)2 + e = 0\) by first solving: \(\Large d = \frac b{2a} =???\) \(\Large e = c - \frac{b^2}{4a} = ??? \)

OpenStudy (iwanttogotostanford):

ok, and i go C

OpenStudy (jack1):

correct, C

OpenStudy (iwanttogotostanford):

it is really? i did it right?:-D

OpenStudy (jack1):

yep

OpenStudy (iwanttogotostanford):

tjanks!

OpenStudy (jack1):

np

OpenStudy (iwanttogotostanford):

could you help with one more? @Jack1

OpenStudy (jack1):

wait... neg 4 negative 4! dammit, sorrt D not C

OpenStudy (jack1):

*sorry ;(

OpenStudy (iwanttogotostanford):

aww ok:-( thanks again! can i ask you one more

OpenStudy (jack1):

sure

OpenStudy (iwanttogotostanford):

Using the completing-the-square method, find the vertex of the function f(x) = –2x^2 + 12x + 5 and indicate whether it is a minimum or a maximum and at what point. Maximum at (–3, 5) Minimum at (–3, 5) Maximum at (3, 23) Minimum at (3, 23)

OpenStudy (jack1):

ok, so have u completed the square?

OpenStudy (iwanttogotostanford):

yes

OpenStudy (jack1):

and ur equation is now...?

OpenStudy (iwanttogotostanford):

2x^2+35x?

OpenStudy (jack1):

no. what are a, b and c in ur original equation?

OpenStudy (iwanttogotostanford):

X2 8 -2

OpenStudy (jack1):

no. \(\large ax^2 + bx + c = 0\) \(\large 2x^2 + 12x + 5 = 0\) so a is 2 b is 12 c is 5... right?

OpenStudy (jack1):

u with me?

OpenStudy (jack1):

r u still there?

OpenStudy (iwanttogotostanford):

yes! sorry what now? @jack1